# DeepRWCap: Neural-Guided Random-Walk Capacitance Solver for IC Design

### v1

@@ 1 Introduction
Random walk methods provide a scalable stochastic framework for solving elliptic partial differential equations (PDEs) such as Poisson and diffusion equations. Originating from the classical Walk-on-Spheres (WoS) algorithm (muller1956some), they leverage the duality between PDEs and stochastic processes. They have gained renewed interest in computer graphics and deep learning applications (sawhney2020monte; li2023neural; miller2024differential; nam2024poisson) due to the inherent parallelism and controllable accuracy. However, while WoS has been extended to support varying diffusion coefficients (sawhney2022grid), it requires the coefficient to be twice-differentiable and is inapplicable to the piecewise constant dielectrics in IC technologies. Thus, the random walk method for capacitance extraction still requires computationally expensive transition evaluations to guide each step when dealing with complex multi-dielectric domains (visvardis2023deep; huang2024enhancing; huang2025efficientfrwtransitionsstochastic).
@@ 4.3 Network Architecture for Transition Kernel Prediction
Given the structural similarity between the sign distribution $\mathbf{s}_{\alpha}$ and the gradient kernel $\mathbf{g}_{\alpha}$, we propose to learn the signed gradient kernel $\mathbf{s}_{\alpha}\mathbf{g}_{\alpha}$ rather than modeling them separately. However, the gradient kernel is equivariant only to a subgroup of cube symmetries, necessitating at least two specialized face predictors to properly characterize the tangent (parallel to Gaussian surface) and normal (perpendicular to Gaussian surface) faces. The normal-face gradient kernels exhibit challenging learning characteristics including sign ambiguities and dual-peaked distributions, motivating the use of deeper network architectures (see Section 5.2). Since the signed target includes both positive and negative values, we do not employ the ReLU activation in the face predictor $\mathcal{G}_{\theta}$ and use MSE as the training loss:
@@ 4.4 High-Throughput Inference Implementation
When a walker encounters a non-stratified transition domain, the system employs the lightweight models described in Section 4.3, following the two-stage pipeline shown in Figure 5. To maximize GPU utilization, transition domain sampling is performed in batches rather than individually. We implement a producer-consumer architecture where each thread manages a walker pool and submits transition tasks to a lock-free queue.
@@ 5.1 Experimental Setup
All experiments were conducted on a server equipped with an Intel Xeon Silver 4214 CPU @ 2.20GHz and an NVIDIA RTX 4090 GPU. The transition cubes were discretized with $N=23$. The TensorRT compilation and custom CUDA kernel optimizations described in Section 4.4 are critical for achieving practical deployment speeds as demonstrated in Figure 6.
@@ 5.2 Model Training
All models were trained using batch size of 16 for 200 epochs with the AdamW optimizer using $\beta_{1}=0.9$, $\beta_{2}=0.999$, and no weight decay. Gradient clipping was applied with a maximum norm of 1.0. The learning rate schedule employed cosine annealing from $1\times 10^{-3}$ to $5\times 10^{-6}$ with a warmup period during the first 20 epochs. For the dual-objective loss function in (9), we used $\lambda=1$ as the weight. Each full training run required 12.3 hours.
@@
The face solver architecture (see Fig. 4) comprises a positional encoding layer followed by a channel projection using $1\times 1$ convolutions (from $23+2$ input channels to $16$ channels), four depthwise separable layers with channel progression $(16,16,8,4,2)$ and increasing dilation rates $(1,1,2,3)$, and a final $1\times 1$ convolution head. The dilated convolutions enable the model to capture larger spatial contexts while maintaining computational efficiency and producing smoother output distributions. Each depthwise separable block consists of a $3\times 3$ depthwise convolution, batch normalization, GELU activation, followed by a $1\times 1$ pointwise convolution, batch normalization, and GELU activation. The tangent gradient solver variant follows the same structure, and the normal gradient solver includes wider channels $(64,64,32,32,16,16,8,4)$. The 3D face selector networks (see Fig. 3) use standard 3D convolutions with a stride of 2, batch normalization, GELU activations, followed by fully connected layers. We use channel progressions with $(1,4,8,16)$ channels for Poisson kernel selection, $(1,8,16,64)$ channels for gradient kernel selection.
Figure 6: Performance analysis after introducing custom CUDA Kernels and TensorRT compilation, benchmarked with a batch size of 2048: (a) Poisson Sampling and (b) Gradient Sampling.
@@ 5.3 Capacitance Extraction Results
Cases 1-6 involve simple parallel-wire structures across different metal layers and technologies. Cases 7-10 are FinFET structures with local interconnects, with cases 9-10 being particularly challenging due to high-contrast conformal dielectrics. For all cases, the FRW process terminates when the estimated stochastic error falls below 1%. Ground truth values were obtained using the commercial solver Raphael, which requires several hours per case.
@@
Table 2 compares method accuracy across test cases, with each method evaluated over 10 independent trials. The percentage of non-stratified domains requiring neural network inference varies significantly: Poisson kernel sampling in 11.2-55.8% of domains and gradient kernel sampling in 0.0-13.4% of domains.
[…]
@@
5.4 Ablation Study
We conduct an ablation study to assess the impact of architectural design choices on accuracy and efficiency for single-face Poisson kernel prediction, focusing on the core prediction models rather than the relatively standard 3D CNN face selector architectures. We evaluate model variants under consistent training settings in Section 5.2. The MLP baseline includes three hidden layers of 2048 units each. We also evaluate with standard 2D convolutions. As a third baseline, we replace the depthwise separable convolution layers with 3D convolutional layers with stride-2 downsampling to reduce depth while preserving width and height.

### v2

@@ 1 Introduction
Random walk methods provide a scalable stochastic framework for solving elliptic partial differential equations (PDEs) such as Poisson and diffusion equations. Originating from the classical Walk-on-Spheres (WoS) algorithm (muller1956some), they leverage the duality between PDEs and stochastic processes. They have gained renewed interest in computer graphics and deep learning applications (sawhney2020monte; li2023neural; miller2024differential; nam2024poisson) due to the inherent parallelism and controllable accuracy. However, while WoS has been extended to support varying diffusion coefficients (sawhney2022grid), it requires the coefficient to be twice-differentiable and is inapplicable to the piecewise constant dielectrics in ICs. Thus, the random walk method for capacitance extraction still requires computationally expensive transition evaluations to guide each step when dealing with complex multi-dielectric domains (visvardis2023deep; huang2024enhancing; huang2025efficientfrwtransitionsstochastic).
@@ 4.3 Network Architecture for Transition Kernel Prediction
Given the structural similarity between the sign distribution $\mathbf{s}_{\alpha}$ and the gradient kernel $\mathbf{g}_{\alpha}$, we propose to learn the signed gradient kernel $\mathbf{s}_{\alpha}\mathbf{g}_{\alpha}$ rather than modeling them separately. However, the gradient kernel is equivariant only to a subgroup of cube symmetries, necessitating at least two specialized face predictors to properly characterize the tangent (parallel to Gaussian surface) and normal (perpendicular to Gaussian surface) faces, as shown in Appendix A. The normal-face gradient kernels exhibit challenging learning characteristics including sign ambiguities and dual-peaked distributions, motivating the use of deeper network architectures (see Section 5.2). Since the signed target includes both positive and negative values, we do not employ the ReLU activation in the face predictor $\mathcal{G}_{\theta}$ and use MSE as the training loss:
@@ 4.4 High-Throughput Inference Implementation
When a walker encounters a non-stratified transition domain, the system employs the lightweight models described in Section 4.3, following the two-stage pipeline shown in Figure 5. To maximize GPU utilization, transition sampling is batched. We implement a producer-consumer architecture where each thread manages a walker pool and submits transition tasks to a lock-free queue. Implementation details and pseudocode are provided in Appendix B.
@@ 5.1 Experimental Setup
All experiments were conducted on a server equipped with an Intel Xeon Silver 4214 CPU @ 2.20GHz and an NVIDIA RTX 4090 GPU. The transition cubes were discretized with $N=23$. The TensorRT compilation and custom CUDA kernel optimizations described in Section 4.4 are critical for achieving practical deployment speeds as shown in Figure 6.
@@ 5.2 Model Training
All models were trained using batch size of 16 for 200 epochs with the AdamW optimizer using $\beta_{1}=0.9$, $\beta_{2}=0.999$, and no weight decay. Gradient clipping was applied with a maximum norm of 1.0. The learning rate schedule employed cosine annealing from $1\times 10^{-3}$ to $5\times 10^{-6}$ with a warmup period during the first 20 epochs. For the dual-objective loss function in (9), we used $\lambda=1$. Each full training run required 12.3 hours.
@@
The face solver architecture (see Fig. 4) comprises a positional encoding layer followed by a channel projection using $1\times 1$ convolutions (from $23+2$ input channels to $16$ channels), four depthwise separable layers with channel progression $(16,16,8,4,2)$ and increasing dilation rates $(1,1,2,3)$, and a final $1\times 1$ convolution head. The dilated convolutions enable the model to capture larger spatial contexts while maintaining computational efficiency and producing smoother output distributions. Each depthwise separable block consists of a $3\times 3$ depthwise convolution, batch normalization, GELU activation, followed by a $1\times 1$ pointwise convolution, batch normalization, and GELU activation. The tangent gradient solver variant follows the same structure, and the normal gradient solver includes wider channels $(64,64,32,32,16,16,8,4)$. The 3D face selector networks (see Fig. 3) use standard 3D convolutions with a stride of 2, batch normalization, GELU activations, followed by fully connected layers. We use channel progressions of $(1,4,8,16)$ for Poisson kernels and $(1,8,16,64)$ for gradient kernels.
Figure 6: Performance analysis after introducing custom CUDA Kernels and TensorRT compilation, benchmarked with a batch size of 2048: (a) Poisson and (b) Gradient.
@@ 5.3 Capacitance Extraction Results
Cases 1-6 involve simple parallel-wire structures across different metal layers and technologies. Cases 7-10 are FinFET structures with local interconnects, with cases 9-10 being particularly challenging due to high-contrast conformal dielectrics. For all cases, the FRW process terminates when the estimated stochastic error falls below 1%. Ground-truth values are obtained from Raphael, which takes hours per case. A detailed breakdown of test case structures and example visualizations is provided in Appendix C.
@@
Table 2 compares method accuracy across test cases, with each method evaluated over 10 independent trials. The percentage of non-stratified domains requiring neural network inference varies significantly: Poisson kernel sampling in 11.2-55.8% of domains and gradient kernel sampling in 0.0-13.4% of domains. Runtime comparisons are shown in Figure 7. Per-case runtimes for all methods, including FRW-FDM and GE-CNN, are listed in Appendix D. Overall, cases 7-10 present shorter runtimes because the smaller conductor spacing accelerates FRW convergence. Statistical analysis using Wilcoxon signed-rank tests ($\alpha=0.05$) shows DeepRWCap achieves a $1.23\times$ speedup over Microwalk ($p=0.024$) while maintaining statistically equivalent accuracy. Although AGF achieved faster execution times, it exhibited poor reliability with high variance in error rates ($5.18\%\pm 7.81\%$ vs DeepRWCap’s $1.24\%\pm 0.53\%$).
5.4 Ablation Study
We conduct an ablation study to assess the impact of architectural design choices on accuracy and efficiency for single-face Poisson kernel prediction. We evaluate model variants under consistent training settings in Section 5.2. The MLP baseline includes three hidden layers of 2048 units each. We also evaluate with standard 2D convolutions. As a third baseline, we replace the depthwise separable convolution layers with 3D convolutional layers with stride-2 downsampling to reduce depth while preserving width and height. Architectural specifications for all ablation models are provided in Appendix E.
@@ Acknowledgments
Appendix A Additional Details for Gradient Kernel Prediction
The face Poisson kernel maintains the same fundamental shape regardless of face orientation, but the gradient kernel exhibits significantly different characteristics depending on the face’s position and orientation, with its structure varying dramatically based on whether we consider faces that are tangential or normal to the gradient direction, as illustrated in Figure 1.
Figure 1: Unfolded Poisson kernel (left) and gradient kernel (right).
For faces oriented tangentially to the gradient direction (such as the top and bottom faces when computing the gradient with respect to the $z$-axis), the gradient kernels are symmetric, differing only in sign. However, for faces oriented normal to the gradient direction (the side faces), the gradient kernel exhibits a fundamentally different structure characterized by two distinct peaks: a positive peak at one end and a negative peak at the other. This bipolar structure makes the prediction task significantly more challenging, as the model must accurately capture both the magnitude and spatial distribution of these opposing peaks while preserving the correct sign relationships.
The increased complexity of gradient kernel prediction for normal-oriented faces is quantified in Table 1. We trained our smallest model configuration using L1 output normalization with the absolute values while preserving the sign on three different tasks: Poisson kernel prediction on face 0, gradient kernel prediction on face 1 (tangential orientation), and gradient kernel prediction on face 2 (normal orientation). The results demonstrate that the prediction loss on the validation set for the normal-oriented face is approximately one order of magnitude higher than for the tangential face and the Poisson kernel task, confirming the fundamental difficulty in modeling the bipolar gradient structure. This substantial increase in prediction difficulty for normal-oriented faces necessitated the development of a larger model architecture.
Kernel MSE Loss Rel. L2 Loss (%)
Poisson $4.20\times 10^{-8}$ $5.64$
Gradient (Tangential) $4.67\times 10^{-8}$ $5.34$
Gradient (Normal) $94\times 10^{-7} 11.$8311.83
Table 1: Gradient kernel prediction loss comparison across different tasks using the smallest model with L1 output normalization (trained with batch size of 256).
Appendix B DeepRWCap Implementation Details
Algorithm 1 Producer: Random Walker Thread
Input: Walker count $N_{w}$, termination criterion
Output: Capacitance matrix $\hat{\mathbf{C}}$
1: Initialize a walkers pool $P$ of $N_{w}$ walkers
2: Initialize samplers $Q_{\text{gradient}}$ and $Q_{\text{poisson}}$
3:while termination criterion is not met do
4:  for each walker $p\in P$ do
5:   if $p.\mathsf{state}=\mathsf{Terminated}$ then
6:    Update $\hat{\mathbf{C}}$ with the weight $\{w_{\alpha},\mathbf{s}_{\alpha}\}$
7:    $p.\mathsf{state}\leftarrow\mathsf{Active}$
8:   else if $p.\mathsf{state}=\mathsf{WaitingForGPU}$ then
9:    if $\mathsf{IsDataAvailable}(p)$ then
10:     if $\mathsf{IsFirstStep}(p)$ then
11:      $\{\mathbf{v}_{\alpha},w_{\alpha}\}\leftarrow\mathsf{RetrieveResult}(p,Q_{\text{gradient}})$
12:      $\mathsf{PerformTransition}(p,\mathbf{v}_{\alpha},w_{\alpha})$
13:     else
14:      $\mathbf{v}_{\alpha}\leftarrow\mathsf{RetrieveResult}(p,Q_{\text{poisson}})$
15:      $\mathsf{PerformTransition}(p,\mathbf{v}_{\alpha})$
16:     end if
17:     if $p$ hits a conductor then
18:      $p.\mathsf{state}\leftarrow\mathsf{Terminated}$
19:     else
20:      $p.\mathsf{state}\leftarrow\mathsf{Active}$
21:     end if
22:    end if
23:   else if $p.\mathsf{state}=\mathsf{Active}$ then
24:    $S\leftarrow\mathsf{GetCurrentTransitionCube}(p)$
25:    if $\mathsf{IsStratified}(S)$ then
26:     $\mathsf{PerformTransitionAGF}(p,S)$
27:     if $p$ hits a conductor then
28:      $p.\mathsf{state}\leftarrow\mathsf{Terminated}$
29:     end if
30:    else
31:     if $\mathsf{IsFirstStep}(p)$ then
32:      $\mathsf{Submit}(\{p,S\},Q_{\text{gradient}})$
33:     else
34:      $\mathsf{Submit}(\{p,S\},Q_{\text{poisson}})$
35:     end if
36:     $p.\mathsf{state}\leftarrow\mathsf{WaitingForGPU}$
37:    end if
38:   end if
39:  end for
40:end while
The DeepRWCap inference process operates through a producer-consumer architecture where random walker threads (producers) generate sampling tasks and sampler threads (consumers) provide transition samples to guide the next step of the walk. The sampler threads handle GPU data transfers, synchronization, and CUDA kernel launches for batched inference. Algorithm 1 details the producer thread execution, with the following key function definitions:
Walker State Management:
  • $\mathsf{IsDataAvailable}(p)$: Checks if sampling results are ready for walker $p$.
  • $\mathsf{IsFirstStep}(p)$: Determines if walker $p$ is at its initial position (requiring gradient kernel).
  • $\mathsf{GetCurrentTransitionCube}(p)$: Extracts the local dielectric configuration around walker $p$.
  • $\mathsf{IsStratified}(S)$: Tests if dielectric cube $S$ contains only stratified materials.
Task Submission and Retrieval:
  • $\mathsf{Submit}(\{p,S\},Q)$: Submits walker $p$ and dielectric configuration $S$ to the appropriate sampler queue $Q$.
  • $\mathsf{RetrieveResult}(p,Q)$: Retrieves processed results from sampler queue $Q$ for walker $p$. Returns $\{\mathbf{v}_{\alpha},w_{\alpha}\}$ (transition vector and weight) for gradient sampling or $\mathbf{v}_{\alpha}$ (transition vector only) for Poisson sampling.
Transition Execution:
  • $\mathsf{PerformTransition}(p,\mathbf{v}_{\alpha},w_{\alpha})$: Updates walker $p$’s position using transition vector $\mathbf{v}_{\alpha}$ and records the weight $w_{\alpha}$ for first steps.
  • $\mathsf{PerformTransition}(p,\mathbf{v}_{\alpha})$: Updates walker $p$’s position using transition vector $\mathbf{v}_{\alpha}$ for subsequent steps.
  • $\mathsf{PerformTransitionAGF}(p,S)$: Executes analytical transition using the AGF approach (huang2024enhancing) for stratified regions.
Queue Management: The algorithm maintains separate queues for the gradient sampler ($Q_{\text{gradient}}$) and the Poisson sampler ($Q_{\text{poisson}}$), enabling specialized batch processing for each kernel type. Walkers transition between $\mathsf{Active}$, $\mathsf{WaitingForGPU}$, and $\mathsf{Terminated}$ states as they progress through the dielectric structure until reaching conductor boundaries.
Appendix C Test Case Description
Table 2 summarizes the structural characteristics of the ten capacitance-extraction test cases, including their layer counts, conformal regions, block decompositions, and conductor counts. Figure 2 provides example visualizations for Cases 7–10, illustrating the variation in geometric complexity across the dataset. These cases highlight the more intricate dielectric and conductor configurations that drive the differences observed in the runtime comparison.
Figure 2: Test case visualizations. (a) Case 7, (b) Case 8, (c) Case 9, and (d) Case 10
Case # Layers # Conformal # Blocks # Conductors
1 58 1482 1888 251
2 57 5838 14592 331
3 58 7062 $83$ 1181
4 50 42 71 24
5 32 38 63 22
6 32 46 75 26
7 57 244 520 12
8 57 150 277 9
9 45 272 914 47
10 42 1114 4511 48
Table 2: Capacitance extraction test cases breakdown.
Appendix D Detailed Runtime Comparison
The detailed runtimes in Table 3 complement the values in Fig. 7. FRW-FDM and GE-CNN runtimes are added for additional context. The high FRW-FDM costs arise from repeatedly solving non-stratified dielectric regions with a finite–difference mesh, making it impractical for large-scale extraction. GE-CNN also shows notable overhead for several practical reasons: (1) its group-equivariant convolutions are custom kernels that cannot benefit from TensorRT, (2) the 24-element symmetry group multiplies the work per convolution, and (3) the GMM sampling procedure does not inherently restrict samples to cube faces, requiring boundary checks and resampling.
Case Runtime (s)
FRW- FDM GE- CNN FRW- AGF Micro- walk DeepRW Cap (ours)
1 2360.7 243.1 9.4 36.6 29.3
2 603.4 38.6 2.3 6.4 9.7
3 1633.7 107.2 13.0 33.1 14.6
4 4323.2 190.1 8.6 33.5 27.4
5 3404.9 88.5 11.7 29.9 22.2
6 3889.0 96.4 8.2 30.2 25.4
7 171.5 16.1 2.9 8.0 3.6
8 97.7 10.7 2.0 7.0 2.2
9 691.8 37.6 11.9 17.8 13.7
10 395.1 25.7 11.3 18.2 9.9
Table 3: Detailed Runtime Comparisons (seconds).
Appendix E Additional Details for Ablation Study
This section provides detailed architectural specifications for the ablation study models.
MLP Baseline
The MLP baseline flattens the $23^{3}$ volumetric input into 12,167 features and processes them through three hidden layers of 2048 units each with BatchNorm1d and GELU activation. The output is reshaped to a $\times23$ face grid. This direct volumetric-to-face mapping approach is the most parameter-heavy baseline due to the large fully connected layers.
3D Convolutional Baseline
[…]
2D Convolutional Baseline
[…]
GE-CNN + GMM
[…]
Depthwise Separable Convolutions
[…]
DS Conv + Learnable Positional Encoding
This variant extends the base depthwise separable architecture by adding two learnable parameter channels (initialized with random normal distribution) to the input, expanding the first layer to 25 channels.
DS Conv + Grid Positional Encoding
Our best-performing architecture concatenates fixed normalized $x,y$ coordinate channels to the input volume, providing explicit spatial awareness. The coordinate grids are registered as non-trainable buffers, with the additional input channels requiring only a small expansion of the first convolutional layer.